The weight of the 5 grade boys is assumed to have normally distributed with mu = 0 kg and sigma = 5 kg find:





The area under the normal curve to the left of point 34kg. What does this area indicate?

1
Expert's answer
2021-11-30T12:46:59-0500

Let "X=" the weight of boy: "X\sim N(\mu, \sigma^2)."

Given "\mu=50\ kg, \sigma=5\ kg." The area under the normal curve to the left of point 34kg indicates the probability that the weight is less than 34 kg


"P(X<34)=P(Z<\dfrac{34-50}{5})"

"=P(Z<-3.2)\approx0.000687"